OpenStudy (anonymous):

the length of a rectangle is increasing as a rate of 8cm/s and the width at 3cm/s. when the length is 20cm and the width is 10cm, how fast is the are of the rectangle increasing?

OpenStudy (saifoo.khan):

area*

OpenStudy (saifoo.khan):

dl/dt = 8 cm/s dw/dt = 3 cm/s A = lw dA/dt = l*dw/dt + w*dl/dt When l = 20 cm and w = 10 cm dA/dt = 20(3) + 10(8) dA/dt = 60 +80 dA/dt = 140 cm²/s

OpenStudy (anonymous):

thanks!